Differentiate the given trigonometric functions.

1. 𝑓(π‘₯) = 3 𝑠𝑖𝑛2π‘₯ π‘π‘œπ‘ 4π‘₯

2. 𝑦 = 1/2 π‘‘π‘Žπ‘›^4(𝑠𝑒𝑐 6π‘₯)

3. 𝑔(𝑑) = tan 𝑑/π‘π‘œπ‘ π‘‘βˆ’4

4. 𝑦 = 8 𝑐𝑠𝑐 π‘₯/4 π‘π‘œπ‘‘ π‘₯/4

5. 𝑓(π‘₯) = π‘₯𝑠𝑖𝑛3π‘₯ + π‘₯ π‘π‘œπ‘ 3π‘₯


1
Expert's answer
2020-12-15T19:51:22-0500

1) Here "f(x)=3.sin2x.cos4x"

So, "f'(x)=\frac{d}{dx}(3.sin2x.cos4x)"

"=3.\frac{d}{dx}(sin3x.cos4x)"

"=3.[sin3x.\frac{d}{dx}(cos4x)+cos4x.\frac{d}{dx}(sin3x)]"

"=3.[sin3x.(-sin4x).\frac{d}{dx}(4x)+cos4x.cos3x.\frac{d}{dx}(3x)]"

"=3.[sin3x.(-sin4x).4+cos4x.cos3x.3]"

"=3.[-4.sin3x.sin4x+3.cos3x.cos4x]"


2) Here "y=(1\/2)tan^4(sec6x)"

Now, "\frac{dy}{dx}=\frac{d}{dx}[\frac{1}{2}.tan^4(sec6x)]"

"=\frac{1}{2}.\frac{d}{dx}[tan^4(sec6x)]"

"=\frac{1}{2}.4.tan^3(sec6x).\frac{d}{dx}[tan(sec6x)]"

"=2.tan^3(sec6x).sec^2(sec6x).\frac{d}{dx}(sec6x)"

"=2.tan^3(sec6x).sec^2(sec6x).(sec6x.tan6x).\frac {d}{dx}(6x)"

"=2.tan^3(sec6x).sec^2(sec6x).sec6x.tan6x.6"

"=12.sec6x.tan6x.sec^2(sec6x).tan^2(sec6x)"


3) Here "g(t)=\frac{tant}{cost-4}"

Therefore, "g'(t)=\frac {d}{dt}[\frac{tant}{cost-4}]"

"=\frac{[(cost-4).\frac{d}{dt}(tant)]-[tant.\frac{d}{dt}(cost-4)]}{(cost-4)^2}"

"=\frac{[(cost-4).sec^2t]-[tant.(-sint)]}{(cost-4)^2}"

"=\frac{sec^2t.(cost-4)+tant.sint}{(cost-4)^2}"


4) Here "y=8.csc(\frac{x}{4}).cot(\frac{x}{4})"

Therefore, "\frac{dy}{dx}=\frac{d}{dx}[8.csc(\frac{x}{4}).cot(\frac{x}{4})]"

"=8[csc(\frac{x}{4}).\frac{d}{dx}(cot(\frac{x}{4}))+cot(\frac{x}{4}).\frac{d}{dx}(csc(\frac{x}{4}))]"

"=8[csc(\frac{x}{4}).(-csc^2(\frac{x}{4})).\frac{d}{dx}(\frac{x}{4})+cot(\frac{x}{4}).(-csc(\frac{x}{4}).cot(\frac{x}{4})).\frac{d}{dx}(\frac{x}{4})]"

"=8[-csc^3(\frac{x}{4}).\frac{1}{4}-csc(\frac{x}{4}).cot^2(\frac{x}{4}).\frac{1}{4}]"

"=-8.\frac{1}{4}.csc(\frac{x}{4})[csc^2(\frac{x}{4})+cot^2(\frac{x}{4})]"

"=-2.csc(\frac{x}{4})[csc^2(\frac{x}{4})+cot^2(\frac{x}{4})]"


5) Here "f(x)=x.sin3x+x.cos3x"

"=x.(sin3x+cos3x)"

Therefore, "f'(x)=\frac{d}{dx}[x.(sin3x+cos3x)]"

"=x.\frac{d}{dx}(sin3x+cos3x)+(sin3x+cos3x).\frac{d}{dx}(x)"

"=x.[cos3x.\frac{d}{dx}(3x)+(-sin3x).\frac{d}{dx}(3x)]+(sin3x+cos3x).1"

"=x.[3.cos3x-3.sin3x]+(sin3x+cos3x)"

"=(1-3x).sin3x+(1+3x).cos3x"